OpenStudy (anonymous):

Simplify 7 - 2(5 - 2x)

OpenStudy (anonymous):

4x-3

OpenStudy (anonymous):

7-2(5-2x) =7-10+4x =-3+4x So, distribute the 2 by doing 2*5 and 2*-2x

OpenStudy (anonymous):

7-10+4x=4x-3

OpenStudy (anonymous):

4X-3 X=3/2

OpenStudy (anonymous):

how it the world does \[x=\frac{3}{2}\]? x is a variable. it could be anything. even my favorite number...